OpenStudy (anonymous):

Multiply and simplify: (3+ radical 6)(5+ radical 6)

OpenStudy (dape):

\[ (3+\sqrt{6})(5+\sqrt{6})=15+3\sqrt{6}+5\sqrt{6}+\sqrt{6}^2=15+8\sqrt{6}+6=21+8\sqrt{6} \]

OpenStudy (anonymous):

simple.

OpenStudy (anonymous):

\[(3+\sqrt{6})(5+\sqrt{6}) = 15+3\sqrt{6}+5\sqrt{6}+6=8\sqrt{6}+21\]